THE ACCURATE DETERMINATION OF MAGNETIC HYPERFINE AND ZEEMAN PARAMETERS FOR 2Π DIATOMIC MOLECULES FROM EXPERIMENTAL DATA

FILIPPO TAMASSIA, Dipartimento di Chimica Fisica ed Inorganica, Facoltà di Chimica Industriale, Università di Bologna, Viale del Risorgimento 4, 40136, Bologna, Italy; JOHN M. BROWN, The Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, OX1 3QZ, UK; \underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada, K1A OR6.

The Hamiltonian for a ^{2}Π state of a diatomic molecule has an indeterminacy between the parameters γ and A_{D}, which can be removed by a unitary transformation to give either an effective γ̃ with Ã_{D} = 0 or an effective Ã_{D} with γ̃ = 0.^{a} The present work extends this transformation to the magnetic hyperfine and Zeeman terms in the Hamiltonian. Among the hyperfine contributions, the term bI·S is particularly affected by the transformation. Care is therefore required in relating the observed value b̃ to the electron spin density. Similar considerations apply to observed Zeeman parameters. Applications to the spectra of NO^{b} and FO^{c} will be presented.

aJ. M. Brown and J. K. G. Watson, J. Mol. Spectrosc. 65, 65 (1977).

bT. D. Varberg, F. Stroh, and K. M. Evenson, J. Mol. Spectrosc. 196, 5 (1999).

cF. Tamassia, J. M. Brown, and K. M. Evenson, J. Chem. Phys. 110, 7273 (1999).